OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (uri):

The slope of y = 4x + 7 is 4 then the slope of the perpendicular is -1/4.

OpenStudy (uri):

And then you have slope and a point,now find y intercept :y =mx +b

OpenStudy (anonymous):

now finding the y intercept. wud m be -1?

OpenStudy (anonymous):

-3=8+x sorry

OpenStudy (anonymous):

thus thus x is -15

OpenStudy (anonymous):

final answer is : y=8-15

OpenStudy (anonymous):

how did u get tht

OpenStudy (anonymous):

i got it by: finding slope which is 4

OpenStudy (anonymous):

then substituting (2,-3) into equation with slope of 4

OpenStudy (anonymous):

so y=4x + something

OpenStudy (anonymous):

-3=4(2)+something

OpenStudy (anonymous):

find the something